# GORAG: Graph-based Retrieval Augmented Generation for Dynamic Few-shot Text Classification

### v3

@@ 4.2.1. Effectiveness Evaluation
As shown in Table 2, GORAG achieves the best classification accuracy over all four rounds with 1-shot and 5-shot labeled indexing data on WOS dataset, it surpasses all RAG-based baselines as well as the state-of-the-art model Entailment. Compared with Entailment, when apply 1-shot setting, GORAG achieves at most 31.6% accuracy gain at the first round. On Reuters dataset, GORAG achieves the best classification accuracy over the last 3 rounds with 1-shot indexing data.
Furthermore, based on the experiment results in Table 2, we can have following observations:
  • Firstly, NN-based Entailment fails on Reuters due to its need for text-formed label names. As noted in prior studies (Liu et al., 2023; Liu et al., 2024c; Lei et al., 2023), it overfits easily, with recall and hallucination issues shown in Table 8. Further analysis is in Appendix subsection E.3.
  • Secondly NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long LLM inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from rounds 2 to 4. While Propositionizer retrieves more fine-grained results, its lack of task awareness results in consistently poor performance on DFSTC tasks. On the Reuters dataset, both Long Context RAG baselines fail when label names are unavailable, as the absence of semantic-rich labels prevents effective classification, especially with long LLM inputs.
  • Thirdly, compression-based RAG models shorten LLM inputs and emphasize labels and key terms but only slightly improve performance across datasets. They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals. Additionally, failing to compress label sets results in lengthy LLM inputs as new labels are introduced in each round.
  • Lastly, GORAG addresses the issues of Compression-based RAG models. From R1 to R4, GORAG adapts better to dynamic label updates. On the Reuters dataset, when label names are not available, GORAG achieves comparable or better performance than other baselines. According to Table 3, a vast amount of knowledge is added to the weighted graph by the online indexing mechanism, and contributes to our effectiveness.
@@ 4.2.2. Efficiency Evaluation
(a) 1-shot setting.
(b) 5-shot setting.
(c) Time cost on WOS dataset.
(d) Time cost on Reuters dataset.
Figure 5. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds, and model Training/Indexing time cost on two datasets under 1-shot setting.
To evaluate the efficiency of GORAG, we compare GORAG’s indexing time cost with other baselines’ training or indexing time cost on the 4 round split version of WOS and Reuters datasets, the results are shown in Figure 5(c) and 5(d). For both datasets, Propositionizer achieves the worst indexing efficiency. This is because it employs an additional fine-tuned LLM to convert each side information into propositions. In this process, this extra LLM rewrite each labeled or query text and generates a long list of propositions, introducing more efficiency overhead compared to other models that use the LLM solely for classification or keyword extraction, which typically generate only a few tokens. It is also worth noting that, although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines.
4.3. Zero-shot Experiments
To further study GORAG’s 0-shot ability, we compare GORAG’s performance with some widely applied open-source LLM models without RAG approaches. We conducted experiments on the WOS dataset by providing only the label names, without any labeled data to GORAG for creating the graph, the result is shown in Table 4. In the 0-shot setting, GORAG first generates label descriptions based on the label names, then extracts keywords from these descriptions, without using any information from labeled texts. Note that the Entailment model cannot be applied in this setting, as they require labeled data for training. For other RAG-based baselines, they are considered equivalent to their backbone LLMs in the 0-shot setting due to a lack of external information sources.
As shown by the results, GORAG achieves comparable performance with other open-source LLM models in the first round and outperforms them in all subsequent rounds. This is because, when the number of target labels is small, the benefit of compressing the target label set is not as pronounced. However, as the number of labels increases, the importance of filtering the label set becomes more significant, allowing GORAG to consistently outperform other open-source LLM models.
Table 4. Zero-shot experiment result on WOS dataset.
Model Round
R1 R2 R3 R4
Qwen2-7B (Yang et al., 2024) 0.3322 0.1790 0.1451 0.1444
Mistral0.3-7B (Jiang et al., 2023) 0.1436 0.0540 0.0336 0.0270
LLaMA3-8B (Touvron et al., 2023) 0.3351 0.1614 0.1161 0.0930
GORAG 0.3305 0.2567 0.2230 0.2102
4.4. Generalization Evaluation
To evaluate the generalization ability of GORAG, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than Reuters dataset, and more complex texts than the IFS-Rel dataset. In Figure 5, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset. This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 5, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
@@ 4.4. Generalization Evaluation
WOS GORAG offline 0.3063 0.2302 0.2455 0.2156
@@
Reuters GORAG offline 0.0000 0.1000 0.0917 0.1032
GORAG unit 0.0000 0.0625 0.1667 0.0795
GORAG keyword 0.0000 0.1000 0.1167 0.1438
GORAG 0.0875 0.1750 0.1667 0.1667
4.5. Ablation Study
To analyze GORAG’s superior performance, we conduct an extensive ablation study in the 1-shot setting, as RAG models show minimal performance changes across shot settings. Specifically, we examine how different GORAG components impact classification performance using its variants. GORAG offline removes the online indexing mechanism. GORAG unit removes the edge weighting mechanism, every edge in this variant is assigned with weight 1. GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, this is because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance. Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals. Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
@@ 5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Text Classification (DFSTC) task. Extensive experiments on three datasets with different characteristics demonstrate the effectiveness of GORAG in classifying texts with only a limited number or even no labeled data. Additionally, GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round. With extensive of ablation studies, we confirm that GORAG’s edge weighting, adaptive retrieval, and online indexing mechanisms contribute to its effectiveness. For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
@@ Appendix A Notations
Due to the space limitation, we put the summary table on important notations here.
@@
$r$ The round of label updation.
@@
$\mathcal{Y}^{r}$, $\mathcal{Y}^{r}_{new}$ All target label and new target labels at round $r$.
$\mathcal{D}^{r}$, $\mathcal{D}^{r}_{new}$ The labeled training text for all seen labels at round $r$.
$\mathcal{T}^{r}$, $\mathcal{T}_{new}^{r}$ All seen texts at round $r$.
@@
$\mathcal{K}_{new}^{r}$ The extracted keyword set for new labels of round $r$.
@@
$\mathcal{V}_{exist}$ The keywords that exist in graph.
$\mathcal{V}_{notexist}$ The keywords do not exist in graph.
@@
Appendix B Complexity Analysis
In this section, we analysis the time and space complexity of GORAG’s graph construction, retrieval and prediction procedure. We denote the maximum number of terms of the input text as $m_{t}$, the number of unique terms of training corpus $\mathcal{D}_{new}^{r}$ as $u$, the LLM’s maximum input token length as $m_{i}$, LLM’s its maximum extraction, generation, and classification token length as $m_{e}$, $m_{g}$, and $m_{c}$, respectively.
B.1. Graph Construction Complexity
Firstly, we analysis the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e}))$; If label names are available, the time complexity of generat label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}))$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$ times, and merging graph requires
$O(|\mathcal{Y}^{r}_{new}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{D}_{new}^{r}|(m_{i}+m_{e})+|\mathcal{Y}^{r}_{new}|(m_{i}+m_{g}+m_{e}+|\mathcal{Y}^{r-1}|)+|\mathcal{V}_{new}^{r}|(|\mathcal{D}_{new}^{r}|m_{t}+u))$.
For the space complexity, the graph is stored with the weighted adjency matrix, hence needs $O(|\mathcal{E}^{r}|)$ space; Storing the training corpus at the $r$-th round would need $O(u|\mathcal{D}_{new}^{r}|)$ space; Storing the representive keywords $\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}$ would need $O(|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$ space. Hence the total space complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{E}^{r}|+u|\mathcal{D}_{new}^{r}|+|\mathcal{K}_{\hat{\mathcal{Y}_{t}^{r}}}|)$.
B.2. Retrieval and Classification Complexity
The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (Mehlhorn, 1988b); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)$; The time complexity of the final classification by LLM is $O(m_{i}+m_{c})$. Hence, the total time complexity of GORAG’s adaptive retrieval and classification can be denoted as
$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}|(|\mathcal{T}_{test}^{r}|m_{t}+u)+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|+m_{i}+m_{c})).$
For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(u|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus. Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
$O(|\mathcal{E}^{r}|+u|\mathcal{T}_{test}^{r}|).$
@@ Appendix C Pseudo Code of GORAG
Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
Theorem D.1.
The Adaptive Candidate Type Generation problem is NP-hard.
Proof.
To demonstrate that the Adaptive Candidate Type Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem. Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem. Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$. Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with |$|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes. Thus, we prove that the adaptive candidate type generation problem is NP-hard. ∎
@@ Appendix D Hardness Proof of The Adaptive Candidate Type Generation problem
0:  $\mathcal{D}_{new}^{r}$: Training text set at the $r$-th round and the provided label $y_{i}\in\mathcal{Y}_{new}^{r}$ for each training text $t_{j}$; $p_{extraction},\ p_{gen}$: The extraction and description generation instruction prompt; $\mathcal{G}^{r-1}$: Weighted graph of the previous round $r-1$.
0:  $\mathcal{G}^{r}$: Constructed weighted graph at the $r$-th round.
@@
2:  for each text $t\in\mathcal{T}_{new}^{r}$ do
3:     $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup LLM(p_{extract},t)\cup y_{t}$
@@
6:     for each label $y_{i}\in\mathcal{Y}_{new}^{r}$ do
7:        Get label description $\mathcal{K}_{y_{i}}=LLM(p_{gen},\mathcal{D}^{r})$.
8:        $\mathcal{V}_{label}^{r}=LLM(p_{extract},\mathcal{K}_{y_{i}})\cup\mathcal{V}_{label}^{r}$
9:     end for
10:  end if
11:  Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
12:  for each node $v\in\mathcal{V}_{new}^{r}$ do
13:     $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
14:     $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
15:  end for
16:  Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
17:  Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
18:  return  $\mathcal{G}^{r}$

### v4

@@ 4.2. Few-shot Experiments
(a) Time cost on WOS dataset.
(b) Time cost on CAD dataset.
Figure 3. Total time cost on 2 datasets under 1-shot setting.
@@ 4.2.1. Effectiveness Evaluation
In Table 2, we apply Qwen2.5-7B backbone for CAD and WOS benchmark. Due to cost and API risk control (failing to respond when detecting dirty words) reasons, we only use the GPT-4o backbone on the IFS-Rel news benchmark. As shown in Table 2, GORAG achieves the best classification accuracy over all dynamic rounds on all benchmarks. Based on the experiment results, we have the following observations:
  • Firstly, with the training data increase (R1 to R4), NN-based Entailment suffers from performance decrease on CAD datasets. This is because the CAD datasets have an unbalanced label distribution, as noted in prior studies (Liu et al., 2023, 2024a; Lei et al., 2023), NN-based models that apply data augmentation techniques are prone to overfit when the ground truth label distribution are unbalanced.
  • Secondly, NaiveRAG performs better than Entailment in round 1 of the WOS and IFS-Rel datasets by avoiding overfitting, but its long inputs hinder its ability to adapt to dynamic label updates, leading to a significant performance drop from R1 to R4.
  • Thirdly, Graph-based RAG baselines only slightly improve performance across datasets. They struggle with dynamic label updates due to issues like uniform indexing, threshold dependency, and narrow sources, leading to inaccurate retrievals.
  • Lastly, from R1 to R4, GORAG adapts better to dynamic label updates, addresses the issues of Graph-based RAG models, and achieve better performance than all tested baselines.
Table 3. Graph incremental trends and memory usage on IFS-Rel dataset round 1 with GPT-4o LLM.
Object 25% 50% 75% 100%
Nodes 1644 3238 (+1394) 4482 (+1244) 5299 (+817)
Edges 1771 3265 (+1494) 4597 (+1332) 5503 (+906)
Memory 13.3 KB 27.6 KB 37.0 KB 46.4 KB
@@ 4.2.2. Efficiency Evaluation
To evaluate the efficiency of GORAG, we compare the total time cost of GORAG and other baselines on the 4 round split version of WOS and CAD datasets with Qwen2.5-7B backbone, the results are shown in Figure 3(a) and 3(b). Although GORAG’s online indexing mechanism introduces a vast amount of knowledge into the weighted graph, we can still achieve comparable or better time efficiency with other baselines. We also list the token usage comparison between GORAG and other baselines in Table 4, and the graph’s memory usage and incremental trends in Table 3, where GORAG consumes fewer tokens than current RAG baselines, and only introduce less than 50KB memory cost compared with solely applying LLMs.
4.3. Ablation Study
To further analyze GORAG’s performance, we conduct an extensive ablation study under 1-shot setting with Qwen2.5-7B backbone, as RAG models show similar performance across shot settings. Specifically, we examine how different GORAG components impact classification performance using its variants. GORAG offline removes the online indexing mechanism. GORAG unit removes the edge weighting mechanism, every edge in this variant is assigned with weight 1. GORAG keyword only uses the keyword extracted from the text to create LLM classification input, rather than the whole text.
Table 4. Token cost on IFS-Rel dataset with 1-shot setting.
Model AVG Token per Query AVG Cost per Query (USD)
GPT-4o 634.6 \$0.005
LongLLMLingua 2771.3 \$0.021
NaiveRAG 4537.4 \$0.034
LightRAG 7866.3 \$0.059
GraphRAG 9296.3 \$0.070
GORAG 557.5 \$0.004
As shown in Table 5, firstly, GORAG offline achieves the worst accuracy, because it lacks a comprehensive retrieval source, making its generated candidate labels inaccurate, hence limits its performance. Secondly, GORAG unit outperforms GORAG offline, but still sub-optimal to GORAG, demonstrates the importance of modeling the varying importance between keywords and each label, which leads to accurate retrievals. Lastly, GORAG keywords demonstrates the important information for text classification are mostly text keywords, as it achieves comparable or better performance than other ablated baselines.
@@ 4.3. Ablation Study
WOS GORAG offline&unit 0.2979 0.2287 0.2645 0.2249
GORAG offline 0.3063 0.2302 0.2455 0.2156
@@
4.4. More Generalization Evaluations
4.4.1. Zero-shot Experiments
To further study GORAG’s 0-shot ability, we compare GORAG’s performance with the original powerful LLM GPT-4o (Hurst et al., 2024) without RAG approaches, and some RAG baselines that are applicable to the 0-shot scenario in Table 6. Due to cost constraints, we conducted experiments on the second largest IFS-Rel dataset by providing only the label names, without any labeled data to GORAG for creating the graph. In the 0-shot setting, we input the labels and their LLM generated descriptions as system prompt for each model, and GORAG would extract keywords from label descriptions, without using any information from labeled texts. Note that the Entailment model cannot be applied in this setting, as they require labeled data for training. For other RAG-based baselines, they are considered equivalent to their base LLMs in the 0-shot setting due to a lack of external information. According to Table 6, GORAG can still achieve better performance than all tested baselines.
4.4.2. Round Generalization Evaluation
To evaluate GORAG’s ability to generalize to a larger number of dynamic rounds, we conducted additional experiments on the 6 and 8 round split versions of the WOS dataset under 1-shot and 5-shot settings, as the WOS dataset contains more testing data and labels than CAD dataset, and more lengthy, complex texts than the IFS-Rel dataset. In Figure 4, the results for rounds 1 to 4 are obtained from the 4-round split version, rounds 5 and 6 are obtained from the 6-round split version, and rounds 7 and 8 are obtained from the 8-round split version of the WOS dataset. This setting can ensure a sufficient amount of testing data at each round, As illustrated in Figure 4, the classification accuracy of GraphRAG and NaiveRAG drops significantly from round 1 to 8, while GORAG maintains competitive classification accuracy as the number of rounds increases through 1 to 8, with both 1 and 5-shot setting, demonstrating the importance of a more precise retrieval.
Table 6. 0-shot experiments for GORAG and RAG baselines on IFS-Rel dataset under 0-shot setting.
Model R1 R2 R3 R4
GPT-4o (Hurst et al., 2024) 0.9219 0.5109 0.4063 0.5936
LongLLMLingua (Jiang et al., 2024) 0.9281 0.5297 0.4234 0.5806
LightRAG (Guo et al., 2025) 0.6718 0.3047 0.2135 0.0977
GraphRAG (Edge et al., 2024) 0.5625 0.1875 0.1458 0.0938
GORAG 0.9328 0.5406 0.4375 0.6063
(a) 1-shot setting.
(b) 5-shot setting.
Figure 4. Experiment result of WOS dataset under 1-shot and 5-shot settings for at most 8 rounds.
@@ 5. Conclusion
In this paper, we propose GORAG, a Graph-based Online Retrieval Augmented Generation framework for the Dynamic Few-shot Social media Text Classification (DFSTC) task. Extensive experiments on hate speech and misinformation detection benchmarks with different characteristics demonstrate the effectiveness of GORAG in classifying social media texts with only a limited number or even no labeled data. GORAG shows its effectiveness in adapting to the dynamic updates of target labels by adaptively retrieving candidate labels to filter the large target label set in each update round. For future work, we aim to further enhance GORAG’s performance and explore its application in more general scenarios.
6. Acknowledgements
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, HKUST(GZ) - CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No. P00659, Hong Kong ITC ITF grant PRP/004/22FX, Zhujiang scholar program 2021JC02X170, HKUST-Webank joint research lab.
@@ Appendix A Notations
Due to the space limitations, we put the summary table on important notations here in Table 7.
@@
$r$ The round of label updating.
$\tau$ The keyword extraction threshold.
@@
$\mathcal{Y}^{r}$ New target labels at round $r$.
$\mathcal{D}^{r}$ The labeled training text for all new labels at round $r$.
$\mathcal{T}^{r}$ Unlabeled texts at round $r$.
@@
$\mathcal{K}^{r}$ The extracted keyword set for new labels of round $r$.
@@
$\mathcal{V}^{t}_{exist}$ The keywords extracted from $t$ and exist in graph.
$\mathcal{V}^{t}_{notexist}$ The keywords extracted from $t$ and not exist in graph.
@@
Appendix B Proof
Here we prove our Adaptive Candidate Label Generation problem can be reduced from the NP-hard Steiner Tree problem (Su et al., 2020).
Theorem B.1.
The Adaptive Candidate Label Generation problem is NP-hard.
Proof.
To demonstrate that the Adaptive Candidate Label Generation problem is NP-hard, we provide a simple reduction of our problem from the Steiner Tree problem. Since the Steiner Tree problem is already proven to be NP-hard (Su et al., 2020), we show that there is a solution for the Steiner Tree problem if and only if there is a solution for our problem. Firstly, given a solution $S$ for our problem, we can construct a Steiner Tree by generating a minimum spanning tree for all nodes in $S$ on graph $\mathcal{G}^{r}$ then connecting all nodes from $\mathcal{V}_{exist}^{t}$ to their closest neighbor nodes in $S$. Secondly, any Steiner Tree $ST$ that contains any node $y\in\mathcal{V}^{r}_{t}$ is also a solution of our problem with $|\{y\in\mathcal{V}^{r}_{t}\}|$ nodes. Thus, we prove that the adaptive candidate type generation problem is NP-hard. ∎
@@ Appendix C Pseudo Code of GORAG
Appendix D Complexity Analysis
Firstly, we analyze the time and space complexity of GORAG’s graph construction. For the text keyword extraction procedure, the time complexity would be $O(|\mathcal{D}^{r}|)$; If label names are available, the time complexity of generating label descriptions and extract label keywords would be $O(|\mathcal{Y}^{r}|)$; Calculating the TFIDF and indexing edges to graph would require $O(|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$ times, and merging graph requires $O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|)$. Hence, the total time complexity of GORAG’s graph construction at the $r$-th round would be:
$O(|\mathcal{Y}^{r}||\mathcal{Y}^{r-1}|+|\mathcal{V}_{new}^{r}||\mathcal{D}^{r}|)$.
[…]
The time complexity of GORAG’s adaptive candidate type generation algorithm is the same with the Mehlhorn algorithm, which is $O(|\mathcal{E}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)$ (MEHLHORN1988125); The time complexity of the online indexing mechanism would cost $|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|$; Hence, the total time complexity of GORAG’s adaptive retrieval can be denoted as
$O(|\mathcal{T}_{test}^{r}|(|\mathcal{E}^{r}|+|\mathcal{V}_{notexist}^{r}||\mathcal{T}_{test}^{r}|+|\mathcal{V}^{r}|log|\mathcal{V}^{r}|)).$
For the space complexity, $O(|\mathcal{E}^{r}|)$ to store the graph, and $O(|\mathcal{T}_{test}^{r}|)$ space is needed to store the testing corpus. Hence, the total space complexity of GORAG’s Retrieval and Classification procedure is
$O(|\mathcal{E}^{r}|+|\mathcal{T}_{test}^{r}|).$
@@ Appendix D Complexity Analysis
[…]
@@
2:for each text $t\in\mathcal{T}^{r}$ do
3:  $\mathcal{V}_{text}^{r}=\mathcal{V}_{text}^{r}\cup KeywordExtract(t)\cup y_{t}$
@@
6:  for each label $y_{i}\in\mathcal{Y}^{r}$ do
7:   Get some label related keywords $\mathcal{V}_{label}^{r}=LLM(p_{gen},\mathcal{D}^{r})\cup\mathcal{V}_{label}^{r}$.
8:  end for
9:end if
10: Let $\mathcal{V}_{new}^{r}=\mathcal{V}_{text}^{r}\cup\mathcal{Y}^{r}_{new}\cup\mathcal{V}_{label}^{r}$ be the new node set at the $r$-th round.
11:for each node $v\in\mathcal{V}_{new}^{r}$ do
12:  $\mathcal{E}_{new}^{r}=\mathcal{E}_{new}^{r}\cup e_{v,y_{i}}$
13:  $\mathcal{W}_{new}^{r}=\mathcal{W}_{new}^{r}\cup w_{v,y_{i}}^{r}$
14:end for
15: Let $\mathcal{G}^{r}_{new}=(\mathcal{V}_{new}^{r},\mathcal{E}_{new}^{r})$ be the newly constructed graph for at the $r$-th round.
16: Let $\mathcal{G}^{r}=Merge(\mathcal{G}^{r}_{new},\mathcal{G}^{r-1})$ be the final constructed graph at the $r$-th round.
17:return $\mathcal{G}^{r}$